# HyViLM: Enhancing Fine-Grained Recognition with a Hybrid Encoder for Vision-Language Models

### v2

@@ Abstract
Recently, there has been growing interest in the capability of multimodal large language models (MLLMs) to process high-resolution images. A common approach currently involves dynamically cropping the original high-resolution image into smaller sub-images, which are then fed into a vision encoder that was pre-trained on lower-resolution images. However, this cropping approach often truncates objects and connected areas in the original image, causing semantic breaks. To address this limitation, we introduce HyViLM, designed to process images of any resolution while retaining the overall context during encoding. Specifically, we: (i) Design a new visual encoder called Hybrid Encoder that not only encodes individual sub-images but also interacts with detailed global visual features, significantly improving the model’s ability to encode high-resolution images. (ii) Propose an optimal feature fusion strategy for the dynamic cropping approach, effectively leveraging information from different layers of the vision encoder. Compared with the state-of-the-art MLLMs under the same setting, our HyViLM outperforms existing MLLMs in nine out of ten tasks. Specifically, HyViLM achieves a 9.6% improvement in performance on the TextVQA task and a 6.9% enhancement on the DocVQA task.
@@ 1 Introduction
Figure 1: (a) An overall performance comparison between HyViLM and other existing MLLMs. (b) and (c) are examples of image fragments. The solid red line in the figures represents the cropping boundary. LLaVA-NeXT provides correct answers when pivotal information remains within the cropping boundaries. However, even slight overlaps with the boundaries cause it to misidentify the information.
@@
Inspired by the remarkable achievements of large language models (LLMs)[50, 3, 42], the development of multimodal large language models (MLLMs)[1, 2, 7, 21, 54, 25] is advancing at a rapid pace. Researchers have directed substantial efforts towards broadening the capabilities of LLMs to encompass additional modalities, resulting in significant breakthroughs in assorted vision-language tasks. Nevertheless, the challenge of fine-grained visual recognition is not adequately addressed. This shortcoming arises partly from the constraints inherent in pre-trained visual encoders[49, 37], which often require MLLMs to process images at relatively low resolutions, typically around $224\times 224$[54, 47] or $336\times 336$ [26]. This limitation hinders the precise identification of fine details and contextual elements within images, resulting in less accurate vision understanding and reasoning.
@@
To solve this limitation, recent endeavors are directed toward enabling MLLMs to handle images of high resolution. One approach involves directly utilizing a visual encoder designed to process higher-resolution images. However, this approach necessitates the training of a high-quality visual encoder, which requires significant computational resources, including high-quality, high-resolution training datasets. An alternative effective strategy is dynamic cropping[46, 23, 45, 15, 29, 26], where the high-resolution image is partitioned into a set of lower-resolution sub-images using adaptive algorithms. These segmented sub-images are subsequently processed by a visual encoder that has been previously trained with lower-resolution image data, thereby augmenting the encoder’s ability to capture finer details while ensuring that maintaining computational efficiency. Nevertheless, this cropping method can result in image fragmentation, thereby altering the original spatial context and disrupting positional associations within the image. This becomes particularly problematic when essential information is located at the edges of these cropped segments, as it poses challenges to encoding such information accurately. For instance, as depicted in Figure 1 when the pivotal information does not intersect with the cropping boundaries, the model LLaVA-NeXT[26] provides the correct answer. However, with a slight alteration in the image such that the pivotal information intersects with the cropping boundaries, LLaVA-NeXT fails to identify it correctly.
@@
To alleviate this effect, in this paper, we propose HyViLM enhancing fine-grained recognition. First, we design a visual encoder, Hybrid Encoder, under the dynamic cropping strategy to address the issue of image fragmentation. Hybrid Encoder can acquire comprehensive global information to enhance the encoding of each sub-image. Furthermore, we propose a different way of visual feature interaction called ConvNeXt-ViT Deep Fusion Module (CVFM), which fully utilizes features at different levels for mutual complementarity, achieving optimal visual encoding.
@@
Figure 2 demonstrates the difference between our approach and previous methods. Previous methods include Dynamic Resolution, Mini-Gemini[22], and Channel-wise Concat. Dynamic Resolution simply involves cropping the image and feeding it into CLIP-ViT[10]. Mini-Gemini interacts the low-resolution CLIP-ViT features with a high-resolution auxiliary branch at the last layer, mainly using Cross Attention, while some methods utilize Channel-wise Concat. The innovation of our method lies in our interaction between two high-resolution branches and the novel idea of deep fusion, which successfully addresses the issue of image fragmentation.
@@
To verify the effectiveness of our proposed method, we conducted extensive experiments on ten vision-language tasks, some of which include fine-grained recognition of high-resolution images, such as TextVQA[40] and DocVQA[36]. The experimental results show that our method outperforms existing MLLMs on 9 out of 10 Vision-Language tasks. Specifically, it surpasses LLaVA-NeXT by 9.6% on TextVQA and by 6.9% on DocVQA.
@@
  • Based on our strategy, we developed the HyViLM model, a powerful MLLM that outperforms existing models on multiple benchmarks, demonstrating the effectiveness of our approach.
@@ 2.1 Large Vision-Language Model
Driven by the tremendous success of large language models (LLMs), there has been growing interest in building end-to-end multimodal large language models. Specifically, most existing MLLMs adopt a modular structure, utilizing an intermediate network to project visual features into the word embedding space of an LLM. Then, the LLM completes various vision-language (VL) tasks in an autoregressive manner. The intermediate networks of existing MLLMs can generally be divided into two types: (i) learned queries, such as perceiver resamplers[1] or Q-Formers[21], which use fixed queries to capture features through cross-attention; and (ii) MLP modules, like those in the LLaVA series[25, 24]. The vision encoder functions as the “eye” of the model, allowing it to interpret and examine visual information. This encoder can incorporate different architectures, like the Vision Transformer (ViT)[43] or ConvNeXt[30] which have been used in image classification and object detection[52]. A majority of MLLMs use CLIP-ViT as the vision encoder, which is trained on large image-text datasets, to proficiently extract visual features. However, CLIP-ViT is pretrained on low-resolution images and cannot directly handle high-resolution images, which limits the model’s performance on fine-grained tasks.
@@ 2.2 High-resolution MLLMS
Directly inputting high-resolution images into visual encoders leads to high computational costs, primarily due to the quadratic complexity associated with the Transformer architecture and the substantial increase in the number of visual tokens. To alleviate this issue, existing high-resolution MLLMs can be divided into two main types. One type adopts a dynamic cropping method, cutting images into patches that are then separately input into Vision Transformers (ViTs) trained on low-resolution images. Although this method is simple and efficient, it can lead to the problem of image fragmentation, altering the original context, and disrupting positional relationships. In particular, if key information is located at the boundaries of the cropped images, it becomes difficult to encode it accurately. The other[44, 51, 14] type uses a dual-encoder approach, introducing a high-resolution branch to supplement the low-resolution branch. For example, Vary[44] and Deepseek-VL[31] utilize the Segment Anything Model (SAM)[19] in the high-resolution visual encoder to better capture detailed information, while MiniGemini[22] and LLaVA-HR[33] adopt ConvNeXt. However, the main visual encoder in these models is still CLIP-ViT, whose capabilities are limited by its low resolution. In our work, we designed a visual encoder called Hybrid Encoder specifically for the dynamic cropping strategy. Hybrid Encoder expands the ability of CLIP-ViT to process high-resolution images, solving the problem of image fragmentation and improving the model’s capacity to handle high-resolution images.
@@ 3 Methodology
Figure 3 provides an overview of HyViLM. Initially, the image is dynamically segmented, resulting in several sub-images, including the original low-resolution image. These sub-images are processed by our newly designed visual encoder, the Hybrid Encoder. After passing through a projection layer, the encoded visual features are integrated with the textual features and subsequently fed into the large language model (LLM). In Section 3.1 we briefly introduce the overall model architecture. Section 3.2 presents our newly designed visual encoder, the Hybrid Encoder. In Section 3.3, we describe our optimal fusion strategy, and Section 3.4 outlines the training paradigm for the HyViLM model.
@@ 3.1 Model Architecture
We employ a dynamic segmentation strategy to partition the input image into smaller blocks while preserving the original aspect ratio. A set of predefined high resolutions is established as follows: $336\times 672$, $672\times 336$, $672\times 672$, $1008\times 336$, and $336\times 1008$. The original image $I_{l}$ is resized and padded to the most suitable predefined resolution, ensuring the aspect ratio is maintained, and subsequently segmented into $n_{w}\times n_{h}$ patches, each of size $336\times 336$.
To determine the optimal resolution, we compute the effective resolution ($Res_{eff}$) and the wasted resolution ($Res_{wasted}$) of the input image relative to each predefined resolution. The selection criterion prioritizes maximizing $Res_{eff}$ and, in cases where multiple resolutions yield identical $Res_{eff}$ values, minimizing $Res_{wasted}$. The computations are defined as follows:
@@
Here, $W_{h}$ and $H_{h}$ represent the dimensions of the predefined high resolutions, while $W_{l}$ and $H_{l}$ denote the width and height of the input image, respectively. The resizing factor Scale is calculated to determine the scaling proportion for the input image. This facilitates the derivation of the resized dimensions $W_{s}$ and $H_{s}$. Subsequently, $Res_{eff}$ is computed to ensure it does not surpass the original image’s resolution, and $Res_{wasted}$ quantifies the unused resolution space. For each predefined resolution, the corresponding $Res_{eff}$ and $Res_{wasted}$ are evaluated. The predefined resolution that yields the highest $Res_{eff}$ is selected. In scenarios where multiple resolutions produce the same $Res_{eff}$, the resolution with the lowest $Res_{wasted}$ is chosen. The image is ultimately resized to the selected predefined resolution $(W_{h},H_{h})$, maintaining the aspect ratio, with any residual areas being padded as necessary.
In addition to the segmented patches, a global view of each input image is generated by resizing the image to a fixed size of $336\times 336$, thereby providing a comprehensive overview. Both the global view and the segmented sub-images are processed through a designed Hybrid Encoder to extract image features. These image features are subsequently aligned via a Multi-Layer Perceptron (MLP) and concatenated with corresponding text features. The combined features are then fed into a language decoder to generate the final predictions.
@@ 3.4 Training Paradigm
As shown in Figure 3, we adopted a two-stage approach in training HyViLM, which includes low-resolution pre-training and high-resolution visual instruction fine-tuning. The goal is to align visual features with the language model during the pre-training stage, followed by fine-tuning the language model in the instruction tuning stage. Specifically, during the pre-training phase, we do not use a dynamic cropping strategy; instead, we directly resize the images to 336 $\times$ 336 for $I_{l}$ and 768 $\times$ 768 for $I_{h}$. We enable CVFM and the visual projector to train while keeping the other layers of the language model and Hybrid Encoder frozen. In the instruction fine-tuning stage, we employ a dynamic cropping strategy, and all parts of the model are trainable, including the ConvNeXt used for generating high-resolution features.
@@ 4.1 Implementation
Table 2: Quantitative results on 10 popular benchmarks, including 5 general benchmarks and 5 document understanding-related benchmarks. “MaxRes” means the maximum resolution supported. The meaning of “672(1536)” is to integrate features of an image with a resolution of $1536\times 1536$ into features of an image with a resolution of $672\times 672$. * means that the metric is provided by OpenCompass[8].
Datasets To achieve effective cross-modality alignment and instruction finetuning, we have gathered high-quality datasets from publicly accessible sources. Initially, we focus on pretraining the projector and interaction layers by amassing approximately 1.2M image captions. This includes 558K image-caption pairs extracted from the LLaVA-filtered CC3M dataset[4] and 695K GPT-4V-generated captions from the ALLaVA dataset[5]. For instruction finetuning, diverse datasets are utilized: 643K single- and multi-turn conversations (minus 21K TextCaps data) from the LLaVA[24] dataset, in addition to 100K QA pairs from ShareGPT4V[6], 10K LAION-GPT-4V captions[11], 6K text-only multi-turn dialogues from LIMA[53] and OpenAssistant2[20], and 700K GPT-4V-constructed instruction pairs from the ALLaVA dataset. To enhance OCR-related competencies, we further incorporate 28K QA pairs comprised of 10K DocVQA[36], 4K ChartQA[34], 10K DVQA[17], and 4K AI2D[18] data. Overall, our datasets support about 1.5M instruction-related conversations aimed at improving image comprehension.
Training Settings In this study, we use the ViT-L pretrained by CLIP[37] and ConvNeXt-L pretrained by LAION[39] as the high-resolution visual encoder and the main architecture of Hybrid Encoder, respectively. For the language model, we utilize LLaMA3-8B-Instruct[42]. We optimize the model for 1 epoch using the AdamW optimizer with a cosine learning rate schedule and a warmup ratio of 0.03. During the pretraining stage, we use a peak learning rate of 1e-3 and a batch size of 256, training the projection and interaction layers. In the instruction fine-tuning phase, our main peak learning rate is 1e-5 with a batch size of 64. The learning rate for the visual encoder is reduced to 2e-6 to ensure stability, while the interaction layer remains at 1e-5. We employ the DeepSpeed Zero[38] 2 strategy, completing the optimization in a total of 32 hours on 32 $\times$ A800 GPUs.
4.2 Evaluations
Benchmarks Following prior work, we evaluated our HyViLM on ten benchmarks. These include five general multimodal understanding benchmarks: MMBench[27], MMMU[48], MME[12], MathVista[32], and HallusionBench[13], as well as five document understanding-related benchmarks: TextVQA[40], DocVQA[41], AI2D[18], InfoVQA[35], and OCRBench[28]. This selection was made to demonstrate that our model not only possesses strong fine-grained recognition abilities but also exhibits excellent general capabilities.
MLLMs We compared our model with some state-of-the-art multi-language large models (MLLMs). (1) Normal Resolution MLLMs, such as LLaVA1.5[25], InstructBLIP[9], Qwen-VL[2], and mPLUG-Owl2[47]. (2) High Resolution MLLMs, such as MiniGemini[22], CogAgent[14], Monkey[23], TextMonkey[29], HiRes-LLaVA-336px[16], LLaVA-UHD[45], DocOwl-1.5-Chat[15], DeepSeek-VL[31], MiniGemini-HD[22], and LLaVA-NeXT[26].
4.3 Main Results

### v3

# FILA: Fine-Grained Vision Language Models

@@ Abstract
Recently, there has been growing interest in the capability of multimodal large language models (MLLMs) to process high-resolution images. A common approach currently involves dynamically cropping the original high-resolution image into smaller sub-images, which are then fed into a vision encoder that was pre-trained on lower-resolution images. While this method enables MLLMs to process high-resolution images within input constraints, it often causes semantic fragmentation by truncating objects and connected areas, leading to degraded model performance, particularly for fine-grained details and irregularly shaped objects. To address this limitation, we propose FILA (Fine-Grained Vision Language Model), a novel framework designed to process images of high resolution while retaining the overall context during encoding. Specifically, we: (i) Introduce a Hybrid Encoder that not only encodes individual sub-images but also interacts with detailed global visual features, significantly improving the model’s ability to encode high-resolution images. (ii) Propose an effective feature fusion strategy for the dynamic cropping approach, effectively leveraging information from different layers of the vision encoder. Compared with the state-of-the-art MLLMs under the same setting, our FILA outperforms existing MLLMs in nine out of ten tasks. Specifically, FILA achieves a 9.6% improvement in performance on the TextVQA task and a 6.9% enhancement on the DocVQA task.
@@ 1 Introduction
Figure 1: (a) An overall performance comparison between FILA and other existing MLLMs. (b) and (c) are examples of image fragments. The solid red line in the figures represents the cropping boundary. LLaVA-NeXT provides correct answers when pivotal information remains within the cropping boundaries. However, even slight overlaps with the boundaries cause it to misidentify the information.
@@
Inspired by the remarkable achievements of large language models (LLMs) (Zhang et al., 2022; Brown et al., 2020; Touvron et al., 2023), the development of multimodal large language models (MLLMs) (Alayrac et al., 2022; Bai et al., 2023; Chen et al., 2024b; Li et al., 2023; Zhu et al., 2023; Liu et al., 2024a) is advancing at a rapid pace. Researchers have directed substantial efforts towards broadening the capabilities of LLMs to encompass additional modalities, resulting in significant breakthroughs in assorted vision-language tasks. Nevertheless, the challenge of fine-grained visual recognition is not adequately addressed. This shortcoming arises partly from the constraints inherent in pre-trained visual encoders (Zhai et al., 2023; Radford et al., 2021), which often require MLLMs to process images at relatively low resolutions, typically around $224\times 224$ (Zhu et al., 2023; Ye et al., 2024) or $336\times 336$  (Liu et al., 2024b). This limitation hinders the precise identification of fine details and contextual elements within images, resulting in less accurate vision understanding and reasoning.
@@
To solve this limitation, recent endeavors are directed toward enabling MLLMs to handle images of high resolution. One approach involves directly utilizing a visual encoder designed to process higher-resolution images. However, this approach necessitates the training of a high-quality visual encoder, which requires significant computational resources, including high-quality, high-resolution training datasets. An alternative effective strategy is dynamic cropping (Ye et al., 2023; Li et al., 2024b; Xu et al., 2024; Hu et al., 2024; Liu et al., 2024d; b), where the high-resolution image is partitioned into a set of lower-resolution sub-images using adaptive algorithms. These segmented sub-images are subsequently processed by a visual encoder that has been previously trained with lower-resolution image data, thereby augmenting the encoder’s ability to capture finer details while ensuring that maintaining computational efficiency. Nevertheless, this cropping method can result in image fragmentation, thereby altering the original spatial context and disrupting positional associations within the image. This becomes particularly problematic when essential information is located at the edges of these cropped segments, as it poses challenges to encoding such information accurately. For instance, as depicted in Figure 1 when the pivotal information does not intersect with the cropping boundaries, the model LLaVA-NeXT (Liu et al., 2024b) provides the correct answer. However, with a slight alteration in the image such that the pivotal information intersects with the cropping boundaries, LLaVA-NeXT fails to identify it correctly.
@@
To alleviate this effect, in this paper, we propose FILA enhancing fine-grained recognition. First, we design a visual encoder, Hybrid Encoder, under the dynamic cropping strategy to address the issue of image fragmentation. Hybrid Encoder can acquire comprehensive global information to enhance the encoding of each sub-image. Furthermore, we propose a different way of visual feature interaction called ConvNeXt-ViT Deep Fusion Module (CVFM), which fully utilizes features at different levels for mutual complementarity, achieving optimal visual encoding.
@@
Figure 2 demonstrates the difference between our approach and previous methods. Previous methods include Dynamic Resolution, Mini-Gemini (Li et al., 2024a), and Channel-wise Concat. Dynamic Resolution simply involves cropping the image and feeding it into CLIP-ViT (Dosovitskiy, 2020). Mini-Gemini interacts the low-resolution CLIP-ViT features with a high-resolution auxiliary branch at the last layer, mainly using Cross Attention, while some methods utilize Channel-wise Concat. The innovation of our method lies in our interaction between two high-resolution branches and the novel idea of deep fusion, which successfully addresses the issue of image fragmentation.
@@
To verify the effectiveness of our proposed method, we conducted extensive experiments on ten vision-language tasks, some of which include fine-grained recognition of high-resolution images, such as TextVQA (Singh et al., 2019) and DocVQA (Mathew et al., 2021b). The experimental results show that our method outperforms existing MLLMs on 9 out of 10 Vision-Language tasks. Specifically, it surpasses LLaVA-NeXT by 9.6% on TextVQA and by 6.9% on DocVQA.
@@
  • Based on our strategy, we developed the FILA model, a powerful MLLM that outperforms existing models on multiple benchmarks, demonstrating the effectiveness of our approach.
@@ 2.1 Large Vision-Language Model
Driven by the tremendous success of large language models (LLMs), there has been growing interest in building end-to-end multimodal large language models. Specifically, most existing MLLMs adopt a modular structure, utilizing an intermediate network to project visual features into the word embedding space of an LLM. Then, the LLM completes various vision-language (VL) tasks in an autoregressive manner. The intermediate networks of existing MLLMs can generally be divided into two types: (i) learned queries, such as perceiver resamplers (Alayrac et al., 2022) or Q-Formers (Li et al., 2023), which use fixed queries to capture features through cross-attention; and (ii) MLP modules, like those in the LLaVA series (Liu et al., 2024a; 2023a). The vision encoder functions as the “eye” of the model, allowing it to interpret and examine visual information. This encoder can incorporate different architectures, like the Vision Transformer (ViT) (Vaswani, 2017) or ConvNeXt (Liu et al., 2022) which have been used in image classification and object detection (Zheng et al., 2023). A majority of MLLMs use CLIP-ViT as the vision encoder, which is trained on large image-text datasets, to proficiently extract visual features. However, CLIP-ViT is pretrained on low-resolution images and cannot directly handle high-resolution images, which limits the model’s performance on fine-grained tasks.
@@ 2.2 High-resolution MLLMS
Directly inputting high-resolution images into visual encoders leads to high computational costs, primarily due to the quadratic complexity associated with the Transformer architecture and the substantial increase in the number of visual tokens. To alleviate this issue, existing high-resolution MLLMs can be divided into two main types. One type adopts a dynamic cropping method, cutting images into patches that are then separately input into Vision Transformers (ViTs) trained on low-resolution images. Although this method is simple and efficient, it can lead to the problem of image fragmentation, altering the original context, and disrupting positional relationships. In particular, if key information is located at the boundaries of the cropped images, it becomes difficult to encode it accurately. The other (Wei et al., 2023; Zhang et al., 2024; Hong et al., 2023) type uses a dual-encoder approach, introducing a high-resolution branch to supplement the low-resolution branch. For example, Vary (Wei et al., 2023) and Deepseek-VL (Lu et al., 2024a) utilize the Segment Anything Model (SAM) (Kirillov et al., 2023) in the high-resolution visual encoder to better capture detailed information, while MiniGemini (Li et al., 2024a) and LLaVA-HR (Luo et al., 2024) adopt ConvNeXt. However, the main visual encoder in these models is still CLIP-ViT, whose capabilities are limited by its low resolution. In our work, we designed a visual encoder called Hybrid Encoder specifically for the dynamic cropping strategy. Hybrid Encoder expands the ability of CLIP-ViT to process high-resolution images, solving the problem of image fragmentation and improving the model’s capacity to handle high-resolution images.
@@ 3 Methodology
Figure 3 provides an overview of FILA. Initially, the image is dynamically segmented, resulting in several sub-images, including the original low-resolution image. These sub-images are processed by our newly designed visual encoder, the Hybrid Encoder. After passing through a projection layer, the encoded visual features are integrated with the textual features and subsequently fed into the large language model (LLM). In Section 3.1 we briefly introduce the overall model architecture. Section 3.2 presents our newly designed visual encoder, the Hybrid Encoder. In Section 3.3, we describe our optimal fusion strategy, and Section 3.4 outlines the training paradigm for the FILA model.
@@ 3.1 Model Architecture
We apply a dynamic segmentation method to partition the input image into smaller blocks while maintaining the original aspect ratio. The image is resized and padded to the most appropriate predefined resolution, then segmented into $n_{w}\times n_{h}$ patches, each sized $336\times 336$. The predefined resolutions are: $336\times 672$, $672\times 336$, $672\times 672$, $1008\times 336$, and $336\times 1008$.
To select the best resolution, we calculate the effective resolution ($Res_{eff}$) and the wasted resolution ($Res_{wasted}$) for each predefined resolution. The selection aims to maximize $Res_{eff}$ and minimize $Res_{wasted}$ if multiple resolutions have the same $Res_{eff}$.
@@
Here, $W_{h}$ and $H_{h}$ are the predefined high resolution dimensions, while $W_{l}$ and $H_{l}$ are the input image’s width and height. Scale is used to resize the image proportionally. After resizing, $Res_{eff}$ ensures it does not exceed the input image’s resolution, and $Res_{wasted}$ measures unused space. The predefined resolution with the highest $Res_{eff}$ is chosen, and if multiple resolutions have the same $Res_{eff}$, the one with the least $Res_{wasted}$ is selected.
In addition to the segmented patches, a global view of the image is created by resizing it to $336\times 336$, offering a complete overview. Both the global view and the patches are processed through a Hybrid Encoder to extract features, which are then aligned by an MLP and combined with text features. Finally, these features are input into a language decoder for the final predictions.
@@ 3.4 Training Paradigm
As shown in Figure 3, we adopted a two-stage approach in training FILA, which includes low-resolution pre-training and high-resolution visual instruction fine-tuning. The goal is to align visual features with the language model during the pre-training stage, followed by fine-tuning the language model in the instruction tuning stage. Specifically, during the pre-training phase, we do not use a dynamic cropping strategy; instead, we directly resize the images to 336 $\times$ 336 for $I_{l}$ and 768 $\times$ 768 for $I_{h}$. We enable CVFM and the visual projector to train while keeping the other layers of the language model and Hybrid Encoder frozen. In the instruction fine-tuning stage, we employ a dynamic cropping strategy, and all parts of the model are trainable, including the ConvNeXt used for generating high-resolution features. The detail training setting as described in Section B.
@@ 4.1 Implementation
Datasets To achieve effective cross-modality alignment and instruction finetuning, we have gathered high-quality datasets from publicly accessible sources. Initially, we focus on pretraining the projector and interaction layers by amassing approximately 1.2M image captions. This includes 558K image-caption pairs extracted from the LLaVA-filtered CC3M dataset  (Changpinyo et al., 2021) and 695K GPT-4V-generated captions from the ALLaVA dataset (Chen et al., 2024a). For instruction finetuning, diverse datasets are utilized: 643K single- and multi-turn conversations (minus 21K TextCaps data) from the LLaVA (Liu et al., 2023a) dataset, in addition to 100K QA pairs from ShareGPT4V (Chen et al., 2023), 10K LAION-GPT-4V captions (eV, 2023), 6K text-only multi-turn dialogues from LIMA (Zhou et al., 2023) and OpenAssistant2 (Köpf et al., 2024), and 700K GPT-4V-constructed instruction pairs from the ALLaVA dataset. To enhance OCR-related competencies, we further incorporate 28K QA pairs comprised of 10K DocVQA (Mathew et al., 2021b), 4K ChartQA (Masry et al., 2022), 10K DVQA (Kafle et al., 2018), and 4K AI2D (Kembhavi et al., 2016) data. Overall, our datasets support about 1.5M instruction-related conversations aimed at improving image comprehension.
4.2 Evaluations
Benchmarks Following prior work, we evaluated our HyViLM on ten benchmarks. These include five general multimodal understanding benchmarks: MMBench (Liu et al., 2023b), MMMU (Yue et al., 2024), MME (Fu et al., 2023), MathVista (Lu et al., 2024b), and HallusionBench (Guan et al., 2024), as well as five document understanding-related benchmarks: TextVQA (Singh et al., 2019), DocVQA (Tito et al., 2021), AI2D (Kembhavi et al., 2016), InfoVQA (Mathew et al., 2021a), and OCRBench (Liu et al., 2024c). This selection was made to demonstrate that our model not only possesses strong fine-grained recognition abilities but also exhibits excellent general capabilities.
MLLMs We compared our model with some state-of-the-art multi-language large models (MLLMs). (1) Normal Resolution MLLMs, such as LLaVA1.5 (Liu et al., 2024a), InstructBLIP (Dai et al., 2023), Qwen-VL (Bai et al., 2023), and mPLUG-Owl2 (Ye et al., 2024). (2) High Resolution MLLMs, such as MiniGemini (Li et al., 2024a), CogAgent (Hong et al., 2023), Monkey (Li et al., 2024b), TextMonkey (Liu et al., 2024d), HiRes-LLaVA-336px (Huang et al., 2024), LLaVA-UHD (Xu et al., 2024), DocOwl-1.5-Chat (Hu et al., 2024), DeepSeek-VL (Lu et al., 2024a), MiniGemini-HD (Li et al., 2024a), and LLaVA-NeXT (Liu et al., 2024b).
4.3 Main Results
@@
Table 1: Quantitative results on 10 popular benchmarks, including 5 general benchmarks and 5 document understanding-related benchmarks. “MaxRes” means the maximum resolution supported. The meaning of “672(1536)” is to integrate features of an image with a resolution of $1536\times 1536$ into features of an image with a resolution of $672\times 672$. * means that the metric is provided by OpenCompass (Contributors, 2023).